PMG Agenda August 3, 2000 by fsm30412

VIEWS: 4 PAGES: 13

									D                      PMG Agenda
                      August 3, 2000



 Overview (Weerts)
       Detector status
       Reportable milestones
       Summary
   Silicon (Demarteau)
   Forward Muon (Denisov)
   Installation (Kotcher)
   The next few months (Weerts)




                                       D0 PMG
                                       August 03, 2000
D                         Status Overview


 Silicon
       Marcel will give details
            Production of components looking better
               – Micron is producing, the end is in sight
               – oversight at Promex has improved yield
            Detector looking good
               – Dead/noisy channels reasonable
            Low mass cable rate some concern at
             vendor
               – Production rate has dropped, being
                 monitored carefully; but first half is done
               – Fermilab (Jones group) doing excellent job
                 in completing cables
            10% readout test
               – Continuous struggle, but are reading out
               – F disk will be next




                                                       D0 PMG
                                                       August 03, 2000
D                                 Fiber Tracker Status

 Hookup of clear waveguides
                    Installation to start end of August (
                     platform move completed)
 Waveguide Production and QC are
  going well
                                   Waveguide Production

    350

    300

    250
                                                                      Project Plan
    200
                                                                      Actual
    150

    100

     50

     0
     5/19/00       6/2/00   6/16/00 6/30/00 7/14/00 7/28/00 8/11/00



          60% complete



                                                                       D0 PMG
                                                                       August 03, 2000
D                                    Fiber Tracker Status

 VLPC Cassettes
            36 done
            Sub-component assembly
                  80-98% complete

                  Cassette Production and Characterization

140

120

100

 80

 60

 40

 20

  0
  6-Dec-99   25-Jan-00   15-Mar-00    4-May-00   23-Jun-00   12-Aug-00   1-Oct-00   20-Nov-00   9-Jan-01


                                       Actuals
                                       Schedule Projection
                                       Ready for Delivery to DAB




                                                                                     D0 PMG
                                                                                     August 03, 2000
D                     Status Overview


 Forward Preshower
       Both detectors installed
 ICD
       North & South Installed
 Luminosity monitor
       North & South Installed




                                        D0 PMG
                                        August 03, 2000
D                           Status Overview

 Central Muons
       Commissioning continues
 Forward Muon detectors
       Trig Pixels (48 octants)
            All octants assembled & tested & starting to move
             to DAB
            First A layer plane installed in detector
            Pixel VME crate readout through DAQ
       MDTs (48 octants)
          A layer octants installed

               – Readout through DAQ tomorrow
            B octants: 8 assembled – completion late Aug
            C octants: done & testing
 Gas systems
       Engineering design manpower …. being solved


        More from Denisov




                                                        D0 PMG
                                                        August 03, 2000
D                     Status Overview


 Trigger & DAQ
       Progressing, no schedule concerns
 Online
       No schedule concerns, not on critical path
       Preparing for CFT waveguide installation
        checkout; pushing the DAQ system; need
        full time shifters
       Multiple run functionality going


 Solenoid and beam pipe
       Beam pipe “installed” within D0




                                            D0 PMG
                                            August 03, 2000
D                        Status Overview;
                            Electronics

 Electronics
       Calorimeter
            5000 preamps installed on detector (of
             55k)
            100 pre-production BLS boards in hand
            ECS re-cabled
               – To allow detector move this week
       Muon
            PDT electronics - done
            Scintillator
               – Front end
                   • Have enough in hand, complete early Sept
               – Readout card - done
            MDT - done
            Readout electronics
               – Readout 100%
               – Fanout 15%; complete early Nov, enough in
                 hand for commissioning




                                                        D0 PMG
                                                        August 03, 2000
D                       Status Overview;
                           Electronics

       Tracking Electronics
            Si Interface Card
              – Bids are back; visiting vendor
              – Agreed on final changes to address
                crosstalk issues
              – 2 boards immediately on fast turnaround
              – 15 boards by end Aug
              – Cabling/infrastructure work underway
            Analog Front End boards
              – Delayed to mid-Aug (vendor)
            MCM yields improving
              – 44 good out of 50
              – Unlikely that we need to order additional
                SVX from UTMC “bank” to cover yields
            Adapter card
              – Sept 1: 34 pre-production cards




                                                     D0 PMG
                                                     August 03, 2000
D                           Installation


 Installation
       Lots of activity in DAB
            Muons, Fiber tracker, CPS,
             Calorimeter,Platform moving, detector
             closed ( looks like detector)
       Installation is taking over the upgrade,
        the hall, the experiment, absorbing most
        manpower and Kotcher will talk about it




                                                 D0 PMG
                                                 August 03, 2000
D                             Summary


 Concerns
       There is obvious progress in DØ. Seems
        clear after my absence (is that a sign ? )
       Silicon concerns continue, but shifting
        from production to readout and preparing
        for receipt of Si in DAB
       Electronics delays are a concern for
        commissioning
            Getting the DAB infrastructure in place
       Holding the Nov 1 wall-down date and
        therefore March 2001
            More about that later……..




                                                 D0 PMG
                                                 August 03, 2000
D          The next few months; part 1



    The DØ schedule:

    Up to November 1:
    DØ: Assembles, installs & commissions until
       Nov 1, no access to collision hall, no
       impact on BD ( have lum. monitoring in
       and will try to install SMT radiation
       monitoring)
    BD: brings up complex, establish 36x36
       operations with collsions, measure losses

    Nov 1 thru Feb 28, 01: shutdown of
       Tevatron; DØ continues to assemble,
       rolls in and commissions complete
       detector.




                                             D0 PMG
                                             August 03, 2000
D               The next few months; part 2

    Concerns:    Further problems in Tevatron

    Scenario I:
    CDF (no Si) rolls in in August ( or not), machine
    starts up with no collisions until ????. No
    collisions by Nov 1  pressure to delay Nov 1
    shutdown, because we desire stable conditions
    to operate silicon detectors into beam.
     DØ currently:
    Nov 1 needs to be fixed. If it slips then
    March readiness slips by same amount.

    Scenario II:
     things go so wrong that “we” keep the March 1
    date, but redefine what a Run II detector is.
    Example: no silicon. This decision is made late,
    after we have installed silicon……not good.
    DØ: if this is a conceivable option, we need
    to know by Sept 15, 2000 to delay silicon
    installation

    Other scenarios ??

    We can only conceive of one scenario: our
    default                                 D0 PMG
                                            August 03, 2000

								
To top